Draft 12.24, X3J13/92-102. Thu 9-Apr-1992 2:33am EDT

Programming Language—Common Lisp

8. Structures

Structures i

defstruct Macro

Syntax:

defstruct name-and-options [documentation] {↓slot-description}* → structure-name

name-and-options::=structure-name | (structure-name [[ ↓options ]])
options::=↓conc-name-option |

{↓constructor-option}* | ↓copier-option | ↓include-option | ↓initial-offset-option | ↓named-option | ↓predicate-option | ↓print-function-option | ↓type-option

conc-name-option::=:conc-name | (:conc-name) | (:conc-name conc-name)
constructor-option::=:constructor |

                 (:constructor) |

(:constructor constructor-name) |

(:constructor constructor-name constructor-arglist)

copier-option::=:copier | (:copier) | (:copier copier-name)
predicate-option::=:predicate | (:predicate) | (:predicate predicate-name)
include-option::=(:include included-structure-name {↓slot-description}*)
print-function-option::=(:print-function print-function-name) | (:print-function)
type-option::=(:type type)
named-option::=:named
initial-offset-option::=(:initial-offset initial-offset)
slot-description::=slot-name |

(slot-name [slot-initform [[ ↓slot-option ]]])

slot-option::=:type slot-type |

:read-only slot-read-only-p

Arguments and Values:

conc-name—a symbol name designator. constructor-arglist—a boa lambda list. constructor-name—a symbol.

defstruct

copier-name—a symbol.

included-structure-name—an already-defined structure name. Note that a derived type is not permissible, even if it would expand into a structure name.

initial-offset—a non-negative integer.

predicate-name—a symbol.

print-function-name—a function name or a lambda expression.

slot-initform—a form.

slot-read-only-p—a boolean.

structure-name—a symbol.

type—one of the type specifiers list, vector, or (vector size), or some other type specifier defined by the implementation to be appropriate.

documentation—a string; not evaluated.

Description:

defstruct defines a structured type, named structure-type, with named slots as specified by the slot-options.

defstruct defines readers for the slots and arranges for setf to work properly on such reader functions. Also, unless overridden, it defines a predicate named name-p, defines a constructor function named make-constructor-name, and defines a copier function named copy-constructor-name. All names of automatically created functions might automatically be declared inline (at the discretion of the implementation).

If documentation is supplied, it is attached to structure-name as a documentation string of kind structure, and unless :type is used, the documentation is also attached to structure-name as a documentation string of kind type and as a documentation string to the class object for the class named structure-name.

defstruct defines a constructor function that is used to create instances of the structure created by defstruct. The default name is make-structure-name. A different name can be supplied by giving the name as the argument to the constructor option. nil indicates that no constructor function will be created.

After a new structure type has been defined, instances of that type normally can be created by using the constructor function for the type. A call to a constructor function is of the following form:

(constructor-function-name

 slot-keyword-1 form-1
 slot-keyword-2 form-2

defstruct

…)

The arguments to the constructor function are all keyword arguments. Each slot keyword argument must be a keyword whose name corresponds to the name of a structure slot. All the keywords and forms are evaluated. If a slot is not initialized in this way, it is initialized by evaluating slot-initform in the slot description at the time the constructor function is called. If no slot-initform is supplied, the consequences are undefined if an attempt is later made to read the slot’s value before a value is explicitly assigned.

Each slot-initform supplied for a defstruct component, when used by the constructor function for an otherwise unsupplied component, is re-evaluated on every call to the constructor function. The slot-initform is not evaluated unless it is needed in the creation of a particular structure instance. If it is never needed, there can be no type-mismatch error, even if the type of the slot is specified; no warning should be issued in this case. For example, in the following sequence, only the last call is an error.

 (defstruct person (name 007 :type string))
 (make-person :name "James")
 (make-person)

It is as if the slot-initforms were used as slot-initforms for the keyword parameters of the constructor function.

The symbols which name the slots must not be used by the implementation as the names for the lambda variables in the constructor function, since one or more of those symbols might have been proclaimed special or might be defined as the name of a constant variable. The slot default init forms are evaluated in the lexical environment in which the defstruct form itself appears and in the dynamic environment in which the call to the constructor function appears.

For example, if the form (gensym) were used as an initialization form, either in the constructorfunction call or as the default initialization form in defstruct, then every call to the constructor function would call gensym once to generate a new symbol.

Each slot-description in defstruct can specify zero or more slot-options. A slot-option consists of a pair of a keyword and a value (which is not a form to be evaluated, but the value itself). For example:

 (defstruct ship
   (x-position 0.0 :type short-float)
   (y-position 0.0 :type short-float)
   (x-velocity 0.0 :type short-float)
   (y-velocity 0.0 :type short-float)
   (mass *default-ship-mass* :type short-float :read-only t))

This specifies that each slot always contains a short float, and that the last slot cannot be altered once a ship is constructed.

The available slot-options are:

defstruct

:type type

This specifies that the contents of the slot is always of type type. This is entirely analogous to the declaration of a variable or function; it effectively declares the result type of the reader function. It is implementation-dependent whether the type is checked when initializing a slot or when assigning to it. Type is not evaluated; it must be a valid type specifier.

:read-only x

When x is true, this specifies that this slot cannot be altered; it will always contain the value supplied at construction time. setf will not accept the reader function for this slot. If x is false, this slot-option has no effect. X is not evaluated.

When this option is false or unsupplied, it is implementation-dependent whether the ability to write the slot is implemented by a setf function or a setf expander.

The following keyword options are available for use with defstruct. A defstruct option can be either a keyword or a list of a keyword and arguments for that keyword; specifying the keyword by itself is equivalent to specifying a list consisting of the keyword and no arguments. The syntax for defstruct options differs from the pair syntax used for slot-options. No part of any of these options is evaluated.

:conc-name

This provides for automatic prefixing of names of reader (or access) functions. The default behavior is to begin the names of all the reader functions of a structure with the name of the structure followed by a hyphen.

:conc-name supplies an alternate prefix to be used. If a hyphen is to be used as a separator, it must be supplied as part of the prefix. If :conc-name is nil or no argument is supplied, then no prefix is used; then the names of the reader functions are the same as the slot names. If a non-nil prefix is given, the name of the reader function for each slot is constructed by concatenating that prefix and the name of the slot, and interning the resulting symbol in the package that is current at the time the defstruct form is expanded.

Note that no matter what is supplied for :conc-name, slot keywords that match the slot names with no prefix attached are used with a constructor function. The reader function name is used in conjunction with setf. Here is an example:

 (defstruct (door (:conc-name dr-)) knob-color width material) → DOOR
 (setq my-door (make-door :knob-color 'red :width 5.0))
→ #S(DOOR :KNOB-COLOR RED :WIDTH 5.0 :MATERIAL NIL)
 (dr-width my-door) → 5.0
 (setf (dr-width my-door) 43.7) → 43.7
 (dr-width my-door) → 43.7

defstruct

:constructor

This option takes zero, one, or two arguments. If at least one argument is supplied and the first argument is not nil, then that argument is a symbol which specifies the name of the constructor function. If the argument is not supplied (or if the option itself is not supplied), the name of the constructor is produced by concatenating the string "MAKE-" and the name of the structure, interning the name in whatever package is current at the time defstruct is expanded. If the argument is provided and is nil, no constructor function is defined.

If :constructor is given as (:constructor name arglist), then instead of making a keyword driven constructor function, defstruct defines a “positional” constructor function, taking arguments whose meaning is determined by the argument’s position and possibly by keywords. Arglist is used to describe what the arguments to the constructor will be. In the simplest case something like (:constructor make-foo (a b c)) defines make-foo to be a three-argument constructor function whose arguments are used to initialize the slots named a, b, and c.

Because a constructor of this type operates “By Order of Arguments,” it is sometimes known as a “boa constructor.”

For information on how the arglist for a “boa constructor” is processed, see Section 3.4.6 (Boa Lambda Lists).

It is permissible to use the :constructor option more than once, so that you can define several different constructor functions, each taking different parameters.

defstruct creates the default-named keyword constructor function only if no explicit :constructor options are specified, or if the :constructor option is specified without a name argument.

(:constructor nil) is meaningful only when there are no other :constructor options specified. It prevents defstruct from generating any constructors at all.

Otherwise, defstruct creates a constructor function corresponding to each supplied :constructor option. It is permissible to specify multiple keyword constructor functions as well as multiple “boa constructors”.

:copier

This option takes one argument, a symbol, which specifies the name of the copier function. If the argument is not provided or if the option itself is not provided, the name of the copier is produced by concatenating the string "COPY-" and the name of the structure, interning the name in whatever package is current at the time defstruct is expanded. If the argument is provided and is nil, no copier function is defined.

The automatically defined copier function is a function of one argument. The copier function creates a fresh structure that has the same type as its argument, and that has

defstruct

the same component values as the original structure; that is, the component values are not copied recursively. If the defstruct :type option was not used, the copier function has the same effect as copy-structure, except that additional type-checking might be performed to make sure that the argument to the copier is of type structure-name.

:include

This option is used for building a new structure definition as an extension of another structure definition. For example:

 (defstruct person name age sex)

To make a new structure to represent an astronaut that has the attributes of name, age, and sex, and functions that operate on person structures, astronaut is defined with :include as follows:

 (defstruct (astronaut (:include person)
                       (:conc-name astro-))
    helmet-size
    (favorite-beverage 'tang))

:include causes the structure being defined to have the same slots as the included structure. This is done in such a way that the reader functions for the included structure also work on the structure being defined. In this example, an astronaut therefore has five slots: the three defined in person and the two defined in astronaut itself. The reader functions defined by the person structure can be applied to instances of the astronaut structure, and they work correctly. Moreover, astronaut has its own reader functions for components defined by the person structure. The following examples illustrate the use of astronaut structures:

 (setq x (make-astronaut :name 'buzz
                         :age 45.
                         :sex t
                         :helmet-size 17.5))
 (person-name x) → BUZZ
 (astro-name x) → BUZZ
 (astro-favorite-beverage x) → TANG

 (reduce #'+ astros :key #'person-age) ; obtains the total of the ages
                                       ; of the possibly empty
                                       ; sequence of astros

The difference between the reader functions person-name and astro-name is that person-name can be correctly applied to any person, including an astronaut, while astro-name can be correctly applied only to an astronaut. An implementation might check for incorrect use of reader functions.

defstruct

At most one :include can be supplied in a single defstruct. The argument to :include is required and must be the name of some previously defined structure. If the structure being defined has no :type option, then the included structure must also have had no :type option supplied for it. If the structure being defined has a :type option, then the included structure must have been declared with a :type option specifying the same representation type.

If no :type option is involved, then the structure name of the including structure definition becomes the name of a data type, and therefore a valid type specifier recognizable by typep; it becomes a subtype of the included structure. In the above example, astronaut is a subtype of person; hence

 (typep (make-astronaut) 'person) → true

indicating that all operations on persons also work on astronauts.

The structure using :include can specify default values or slot-options for the included slots different from those the included structure specifies, by giving the :include option as:

(:include included-structure-name {slot-description}*)

Each slot-description must have a slot-name that is the same as that of some slot in the included structure. If a slot-description has no slot-initform, then in the new structure the slot has no initial value. Otherwise its initial value form is replaced by the slot-initform in the slot-description. A normally writable slot can be made read-only. If a slot is read-only in the included structure, then it must also be so in the including structure. If a type is supplied for a slot, it must be a subtype of the type specified in the included structure.

For example, if the default age for an astronaut is 45, then

 (defstruct (astronaut (:include person (age 45)))
    helmet-size
    (favorite-beverage 'tang))

If :include is used with the :type option, then the effect is first to skip over as many representation elements as needed to represent the included structure, then to skip over any additional elements supplied by the :initial-offset option, and then to begin allocation of elements from that point. For example:

 (defstruct (binop (:type list) :named (:initial-offset 2))
   (operator '? :type symbol)
   operand-1
   operand-2) → BINOP
 (defstruct (annotated-binop (:type list)
                             (:initial-offset 3)
                             (:include binop))
  commutative associative identity) → ANNOTATED-BINOP

defstruct

 (make-annotated-binop :operator '*
                       :operand-1 'x
                       :operand-2 5
                       :commutative t
                       :associative t
                       :identity 1)
   → (NIL NIL BINOP * X 5 NIL NIL NIL T T 1)

The first two nil elements stem from the :initial-offset of 2 in the definition of binop. The next four elements contain the structure name and three slots for binop. The next three nil elements stem from the :initial-offset of 3 in the definition of annotated-binop. The last three list elements contain the additional slots for an

annotated-binop.
:initial-offset

:initial-offset instructs defstruct to skip over a certain number of slots before it starts allocating the slots described in the body. This option’s argument is the number of slots defstruct should skip. :initial-offset can be used only if :type is also supplied.

:initial-offset allows slots to be allocated beginning at a representational element other than the first. For example, the form

 (defstruct (binop (:type list) (:initial-offset 2))
   (operator '? :type symbol)
   operand-1
   operand-2) → BINOP

would result in the following behavior for make-binop:

 (make-binop :operator '+ :operand-1 'x :operand-2 5)
→ (NIL NIL + X 5)
 (make-binop :operand-2 4 :operator '*)
→ (NIL NIL * NIL 4)
The selector functions binop-operator, binop-operand-1, and binop-operand-2 would be

essentially equivalent to third, fourth, and fifth, respectively. Similarly, the form

 (defstruct (binop (:type list) :named (:initial-offset 2))
   (operator '? :type symbol)
   operand-1
   operand-2) → BINOP

would result in the following behavior for make-binop:

 (make-binop :operator '+ :operand-1 'x :operand-2 5) → (NIL NIL BINOP + X 5)
 (make-binop :operand-2 4 :operator '*) → (NIL NIL BINOP * NIL 4)

defstruct

The first two nil elements stem from the :initial-offset of 2 in the definition of binop. The next four elements contain the structure name and three slots for binop.

:named

:named specifies that the structure is named. If no :type is supplied, then the structure is always named.

For example:

 (defstruct (binop (:type list))
   (operator '? :type symbol)
   operand-1
   operand-2) → BINOP

This defines a constructor function make-binop and three selector functions, namely

binop-operator, binop-operand-1, and binop-operand-2. (It does not, however, define a

predicate binop-p, for reasons explained below.)

The effect of make-binop is simply to construct a list of length three:

 (make-binop :operator '+ :operand-1 'x :operand-2 5) → (+ X 5)
 (make-binop :operand-2 4 :operator '*) → (* NIL 4)

It is just like the function list except that it takes keyword arguments and performs slot defaulting appropriate to the binop conceptual data type. Similarly, the selector functions

binop-operator, binop-operand-1, and binop-operand-2 are essentially equivalent to car,

cadr, and caddr, respectively. They might not be completely equivalent because, for example, an implementation would be justified in adding error-checking code to ensure that the argument to each selector function is a length-3 list.

binop is a conceptual data type in that it is not made a part of the Common Lisp type system. typep does not recognize binop as a type specifier, and type-of returns list when given a binop structure. There is no way to distinguish a data structure constructed by make-binop from any other list that happens to have the correct structure.

There is not any way to recover the structure name binop from a structure created by make-binop. This can only be done if the structure is named. A named structure has the property that, given an instance of the structure, the structure name (that names the type) can be reliably recovered. For structures defined with no :type option, the structure name actually becomes part of the Common Lisp data-type system. type-of, when applied to such a structure, returns the structure name as the type of the object; typep recognizes the structure name as a valid type specifier.

For structures defined with a :type option, type-of returns a type specifier such as list or (vector t), depending on the type supplied to the :type option. The structure name does not become a valid type specifier. However, if the :named option is also supplied, then the first component of the structure (as created by a defstruct constructor function)

defstruct

always contains the structure name. This allows the structure name to be recovered from an instance of the structure and allows a reasonable predicate for the conceptual type to be defined: the automatically defined name-p predicate for the structure operates by first checking that its argument is of the proper type (list, (vector t), or whatever) and then checking whether the first component contains the appropriate type name.

Consider the binop example shown above, modified only to include the :named option:

 (defstruct (binop (:type list) :named)
   (operator '? :type symbol)
   operand-1
   operand-2) → BINOP

As before, this defines a constructor function make-binop and three selector functions

binop-operator, binop-operand-1, and binop-operand-2. It also defines a predicate

binop-p. The effect of make-binop is now to construct a list of length four:

 (make-binop :operator '+ :operand-1 'x :operand-2 5) → (BINOP + X 5)
 (make-binop :operand-2 4 :operator '*) → (BINOP * NIL 4)

The structure has the same layout as before except that the structure name binop is included as the first list element. The selector functions binop-operator, binop-operand-1, and binop-operand-2 are essentially equivalent to cadr, caddr, and cadddr, respectively. The predicate binop-p is more or less equivalent to this definition:

 (defun binop-p (x)
   (and (consp x) (eq (car x) 'binop))) → BINOP-P

The name binop is still not a valid type specifier recognizable to typep, but at least there is a way of distinguishing binop structures from other similarly defined structures.

:predicate

This option takes one argument, which specifies the name of the type predicate. If the argument is not supplied or if the option itself is not supplied, the name of the predicate is made by concatenating the name of the structure to the string "-P", interning the name in whatever package is current at the time defstruct is expanded. If the argument is provided and is nil, no predicate is defined. A predicate can be defined only if the structure is named; if :type is supplied and :named is not supplied, then :predicate must either be unsupplied or have the value nil.

:print-function

This option can be used only if :type is not supplied. The argument to :print-function should be in a form acceptable to function, and is to be used to print structures of this type. When a structure of this type is to be printed, the argument function is called on three arguments: the structure to be printed, a stream to print to, and an integer indicating the current depth (to be compared against *print-level*).

defstruct

If :print-function and :type are not supplied, or if :print-function is supplied with no argument, then a default printing function is provided for the structure that prints out all its slots using #S syntax. If :print-function and :type are not supplied, but a print function is inherited via the :include option, that print function is used to print the slots.

Supplying :print-function to defstruct is equivalent to defining an appropriate method on the print-object generic function. See Section 7.6.2 (Introduction to Methods).

When *print-circle* is not nil, a user-defined print function can print objects to the supplied stream using write, prin1, princ, or format and expect circularities to be detected and printed using the #n# syntax. This applies to methods on print-object in addition to :print-function options. If a user-defined print function prints to a stream other than the one that was supplied, then circularity detection starts over for that stream. See *print-circle*.

:type

:type explicitly specifies the representation to be used for the structure. Its argument must be one of these types:

vector

This produces the same result as specifying (vector t). The structure is represented as a general vector, storing components as vector elements. The first component is vector element 1 if the structure is :named, and element 0 otherwise.

(vector element-type)

The structure is represented as a (possibly specialized) vector, storing components as vector elements. Every component must be of a type that can be stored in a vector of the type specified. The first component is vector element 1 if the structure is :named, and element 0 otherwise. The structure can be :named only if the type symbol is a subtype of the supplied element-type.

list

The structure is represented as a list. The first component is the cadr if the structure is :named, and the car if it is not :named.

Specifying this option has the effect of forcing a specific representation and of forcing the components to be stored in the order specified in defstruct in corresponding successive elements of the specified representation. It also prevents the structure name from becoming a valid type specifier recognizable by typep.

For example:

 (defstruct (quux (:type list) :named) x y)

defstruct

should make a constructor that builds a list exactly like the one that list produces, with quux as its car.

If this type is defined:

 (deftype quux () '(satisfies quux-p))

then this form

 (typep (make-quux) 'quux)

should return precisely what this one does

 (typep (list 'quux nil nil) 'quux)

If :type is not supplied, the structure is represented as an object of type structure-object.

defstruct without a :type option defines a class with the structure name as its name. The metaclass of structure instances is structure-class.

The consequences of redefining a defstruct structure are undefined.

In the case where no defstruct options have been supplied, the following functions are automatically defined to operate on instances of the new structure:

Predicate

A predicate with the name structure-name-p is defined to test membership in the structure type. The predicate (structure-name-p object) is true if an object is of this type; otherwise it is false. typep can also be used with the name of the new type to test whether an object belongs to the type. Such a function call has the form (typep object 'structure-name).

Component reader functions

Reader functions are defined to read the components of the structure. For each slot name, there is a corresponding reader function with the name structure-name-slot-name. This function reads the contents of that slot. Each reader function takes one argument, which is an instance of the structure type. setf can be used with any of these reader functions to alter the slot contents.

Constructor function

A constructor function with the name make-structure-name is defined. This function creates and returns new instances of the structure type.

Copier function

defstruct

A copier function with the name copy-structure-name is defined. The copier function takes an object of the structure type and creates a new object of the same type that is a copy of the first. The copier function creates a new structure with the same component entries as the original. Corresponding components of the two structure instances are eql.

If a defstruct form appears as a top level form, the compiler must make the structure type name recognized as a valid type name in subsequent declarations (as for deftype) and make the structure slot readers known to setf. In addition, the compiler must save enough information about the structure type so that further defstruct definitions can use :include in a subsequent deftype in the same file to refer to the structure type name. The functions which defstruct generates are not defined in the compile time environment, although the compiler may save enough information about the functions to code subsequent calls inline. The #S reader macro might or might recognize the newly defined structure type name at compile time.

Examples:

An example of a structure definition follows:

(defstruct ship
x-position
y-position
x-velocity
y-velocity
mass)

This declares that every ship is an object with five named components. The evaluation of this form does the following:

1. It defines ship-x-position to be a function of one argument, a ship, that returns the x-position of the ship; ship-y-position and the other components are given similar function definitions. These functions are called the access functions, as they are used to access elements of the structure.

2. ship becomes the name of a type of which instances of ships are elements. ship becomes acceptable to typep, for example; (typep x 'ship) is true if x is a ship and false if x is any object other than a ship.

3. A function named ship-p of one argument is defined; it is a predicate that is true if its argument is a ship and is false otherwise.

4. A function called make-ship is defined that, when invoked, creates a data structure with five components, suitable for use with the access functions. Thus executing

 (setq ship2 (make-ship))

sets ship2 to a newly created ship object. One can supply the initial values of any desired component in the call to make-ship by using keyword arguments in this way:

defstruct

         (setq ship2 (make-ship :mass *default-ship-mass*
                                :x-position 0
                                :y-position 0))

This constructs a new ship and initializes three of its components. This function is called the “constructor function” because it constructs a new structure.

5. A function called copy-ship of one argument is defined that, when given a ship object, creates a new ship object that is a copy of the given one. This function is called the “copier function.”

setf can be used to alter the components of a ship:

 (setf (ship-x-position ship2) 100)

This alters the x-position of ship2 to be 100. This works because defstruct behaves as if it generates an appropriate defsetf for each access function.

;;;
;;; Example 1
;;; define town structure type
;;; area, watertowers, firetrucks, population, elevation are its components
;;;
 (defstruct town
             area
             watertowers
             (firetrucks 1 :type fixnum)    ;an initialized slot
             population
             (elevation 5128 :read-only t)) ;a slot that can't be changed
→ TOWN
;create a town instance
 (setq town1 (make-town :area 0 :watertowers 0)) → #S(TOWN...)
;town's predicate recognizes the new instance
 (town-p town1) → true
;new town's area is as specified by make-town
 (town-area town1) → 0
;new town's elevation has initial value
 (town-elevation town1) → 5128
;setf recognizes reader function
 (setf (town-population town1) 99) → 99
 (town-population town1) → 99
;copier function makes a copy of town1
 (setq town2 (copy-town town1)) → #S(TOWN...)
 (= (town-population town1) (town-population town2)) → true
;since elevation is a read-only slot, its value can be set only
;when the structure is created

defstruct

 (setq town3 (make-town :area 0 :watertowers 3 :elevation 1200))
→ #S(TOWN...)
;;;
;;; Example 2
;;; define clown structure type
;;; this structure uses a nonstandard prefix
;;;
 (defstruct (clown (:conc-name bozo-))
             (nose-color 'red)
             frizzy-hair-p polkadots) → CLOWN
 (setq funny-clown (make-clown)) → #S(CLOWN)
;use non-default reader name
 (bozo-nose-color funny-clown) → RED
 (defstruct (klown (:constructor make-up-klown) ;similar def using other
             (:copier clone-klown)              ;customizing keywords
             (:predicate is-a-bozo-p))
             nose-color frizzy-hair-p polkadots) → klown
;custom constructor now exists
 (fboundp 'make-up-klown) → true
;;;
;;; Example 3
;;; define a vehicle structure type
;;; then define a truck structure type that includes
;;; the vehicle structure
;;;
 (defstruct vehicle name year (diesel t :read-only t)) → VEHICLE
 (defstruct (truck (:include vehicle (year 79)))
             load-limit
             (axles 6)) → TRUCK
 (setq x (make-truck :name 'mac :diesel t :load-limit 17))
→ #S(TRUCK...)
;vehicle readers work on trucks
 (vehicle-name x)
→ MAC
;default taken from :include clause
 (vehicle-year x)
→ 79
 (defstruct (pickup (:include truck))     ;pickup type includes truck
             camper long-bed four-wheel-drive) → PICKUP
 (setq x (make-pickup :name 'king :long-bed t)) → #S(PICKUP...)
;:include default inherited
 (pickup-year x) → 79
;;;
;;; Example 4
;;; use of BOA constructors

;;;
 (defstruct (dfs-boa                      ;BOA constructors
               (:constructor make-dfs-boa (a b c))
               (:constructor create-dfs-boa
                 (a &optional b (c 'cc) &rest d &aux e (f 'ff))))
             a b c d e f) → DFS-BOA
;a, b, and c set by position, and the rest are uninitialized
 (setq x (make-dfs-boa 1 2 3)) → #(DFS-BOA...)
 (dfs-boa-a x) → 1
;a and b set, c and f defaulted
 (setq x (create-dfs-boa 1 2)) → #(DFS-BOA...)
 (dfs-boa-b x) → 2
 (eq (dfs-boa-c x) 'cc) → true
;a, b, and c set, and the rest are collected into d
 (setq x (create-dfs-boa 1 2 3 4 5 6)) → #(DFS-BOA...)
 (dfs-boa-d x) → (4 5 6)

Exceptional Situations:

If any two slot names (whether present directly or inherited by the :include option) are the same under string=, defstruct should signal an error of type program-error.

The consequences are undefined if the included-structure-name does not name a structure type.

See Also:

documentation, print-object, setf, subtypep, type-of, typep, Section 3.2 (Compilation)

Notes:

The argument to :print-function should observe the values of such printer-control variables as *print-escape*.

The restriction against issuing a warning for type mismatches between a slot-initform and the corresponding slot’s :type option is necessary because a slot-initform must be specified in order to specify slot options; in some cases, no suitable default may exist.

The mechanism by which defstruct arranges for slot accessors to be usable with setf is implementation-dependent; for example, it may use setf functions, setf expanders, or some other implementation-dependent mechanism known to that implementation’s code for setf.

copy-structure Function

Syntax:

copy-structure structure → copy

copy-structure

Arguments and Values:

structure—a structure.

copy—a copy of the structure.

Description:

Returns a copy_{6} of the structure.

Only the structure itself is copied; not the values of the slots.

Examples:

See Also:

the :copier option to defstruct

Notes:

The copy is the same as the given structure under equalp, but not under equal.

Table of Contents

8.1 Structures Dictionary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8–1

ii Programming Language—Common Lisp